OpenStudy (anonymous):

What is the value of 4√(4√(4√(4√(4…)) ) ) ?

OpenStudy (anonymous):

\[4\sqrt{4\sqrt{4\sqrt{4\sqrt{4...}}}}\]

OpenStudy (anonymous):

Think it's gonna be 16. The term under the radical converges to 4.

OpenStudy (anonymous):

Total value of the term under the radical, I mean. The actual radicand converges to 16.

OpenStudy (anonymous):

SOLUTION, Pls. or Explanation

OpenStudy (anonymous):